# Fine-tuning Large Language Models for Entity Matching

### v1

@@ 4. Dimension 1: Example Representation
This section investigates how different training example representations affect the performance and generalization capabilities of fine-tuned models. We compare three representations: 1) simple pairs as used in Section 3, 2) pairs augmented by the LLM with textual explanations why the entities match or do not match, and 3) pairs augmented with structured explanations. The complete prompts for instructing the model to generate explanations as well as the augmented training sets are found in the project repository.
@@
Textual Explanations: We evaluate two types of textual explanations generated using GPT-4o-mini. The first approach presents the model with the prompt shown in Figure 2, followed by the label. Afterwards, the model is asked to explain the label. The model is given no guidance on the format of the explanation. This results in detailed explanations, having an average length of 293 tokens, which describe why the entity descriptions match or do not match. As a second approach, we re-run the technique proposed by Wadhwa et al. (Wadhwa et al., 2024), which generates more concise explanations by including short explanations for other entity pairs as demonstrations in the prompt. The resulting explanations have an average length of 90 tokens. Figure 3 shows a training example containing an explanation in the AI part which was generated using this approach. The prompts used for generating these explanations, as well as the longer, more open-ended explanations, are available in the accompanying repository.
Figure 3. Training example containing a textual explanation. The explanation is generated using Wadhwa et al.’s prompt.
@@
Structured Explanations: Our third approach is based on the assumption that LLMs can learn more from structured explanations compared to textual ones. Thus, building on the work of Peeters et al. (Peeters and Bizer, 2024), we instruct the model to generate explanations in a structured format that explicitly mention the attributes that were used for the matching decision, their importance for the decision, as well as the similarity of the attribute values. Figure 4 shows an example of a structured explanation that was generated by GPT-4o-mini. The complete prompt for the generation of the explanation is found in the project repository and is discussed in  (Peeters and Bizer, 2024). The example shows that the LLM incorporates factors beyond string matching for calculating similarities, such as recognizing the relationship between Microsoft and Skype for Business.
Figure 4. Training example including a structured explanation as used during fine-tuning.
@@
Table 3. Fine-tuning results for different types of explanations (F1 scores, deltas to standard fine-tuning on WDC).
Model Train set No Transfer In-Domain Transfer Transfer Cross-Domain Transfer Transfer
@@
4.1. Effectiveness
We compare the effectiveness of different explanation approaches in a non-transfer scenario using the WDC Products dataset. The result of these experiments are found in the No Transfer column of Table 3.
Smaller Models: Llama 8B outperforms standard fine-tuning across all evaluated explanation approaches, as shown in Table 3, column No Transfer. Long textual explanations provide the smallest increase of 1.48 F1 points. The approach by Wadhwa et al. results in a 4.01-point improvement. Structured explanations further improve F1 scores by 0.93 over Wadhwa’s approach. GPT-4o-mini shows a different trend, where only structured explanations outperform the standard fine-tuning, increasing by 0.97 F1 points (Table 3, WDC Products column). The other configurations underperform, with three out of five approaches performing worse than zero-shot (Table 3, Non-transfer column).
Larger Model: Structured explanations are exclusively tested on larger models due to their consistent performance improvement in smaller models. While Llama 70B does not benefit from standard fine-tuning, structured explanations improve its performance by 1.50 F1 points over zero-shot (Table 3, WDC column). In contrast, GPT-4o does not follow this trend. While structured explanations lead to a 1.60-point increase over zero-shot, standard fine-tuning results in a 5.50-point improvement, as shown in the GPT-4o rows in the WDC column of Table 3. Therefore, GPT-4o does not significantly benefit from structured explanations in a non-transfer setting.
4.2. Generalization
@@
In-Domain Generalization: Standard fine-tuning results in Llama 8B achieving 72% of the dedicated models’ performance on product datasets (Table 3, In-Domain Transfer Gain column). The Wadhwa et al. approach and longer textual explanations underperform, achieving 55% and 51%, respectively. However, structured explanations surpass the fine-tuned baseline, with the structured and no importance approaches achieving 93% and 91% of source model performance (Table 3, In-Domain Transfer Gain column).
GPT-4o-mini continues the trend where four out of five explanation approaches underperform the fine-tuned baseline. However, structured explanations outperform the baseline, achieving 23% of the source models’ performance compared to 13% by the fine-tuned baseline (Table 3, In-Domain Transfer Gain column).
Llama 70B benefits from structured explanations, outperforming both the fine-tuned baseline and zero-shot in in-domain generalization. It shows the highest generalization performance on two out of three product datasets (Table 3, In-Domain Transfer Gain column).
As shown in the last row of Table 3, structured explanations do not improve the generalization performance of GPT-4o, as this approach underperforms both the fine-tuned baseline and zero-shot.
Cross-Domain Generalization: Llama 8B struggles to generalize effectively in cross-domain settings, with all approaches underperforming compared to zero-shot performance. In contrast, GPT-4o-mini shows improvements across all explanation approaches, however non-rival zero-shot performance.
In conclusion, structured explanations significantly enhance in-domain generalization, while cross-domain generalization remains limited.
5. Dimension 2: Example Selection and Generation
@@
5.1. Training Set Filtration
The filtration approach is based on the assumption that training data quality outweighs quantity and it is thus beneficial to remove potentially misleading examples from the training set.
Error-based Filtering: GPT-4o-mini is prompted to match pairs in the small WDC training split using the complex-force prompt (see Section 3.3). Training examples for which the model’s labels differ from the ground truth labels are discarded. This approach may discard some correctly labelled examples. However, it is hypothesized that overall dataset quality improves. The resulting training set WDC-filtered has a size of 2006 examples, compared to 2500 examples of the original WDC-small training set (see Table 4).
Relevancy-based Filtering: The second approach involves filtering for relevance, operating on the assumption that fewer, more meaningful examples lead to better outcomes. For example, comparing a hard drive and a TV might pass initial filtration but offer limited value due to their low complexity. By removing such examples, computational demands are certainly reduced, while training results may potentially be enhanced. This approach prompts GPT-4o to select only “interesting” training examples, a term we intentionally leave undefined, allowing the model to interpret it. The model appears to define “interesting” as pairs that share many attributes and are thus highly similar (corner cases). Pairs deemed irrelevant by the model are discarded. The resulting WDC-filtered-rel training set has a total size of 608 examples (see Table 4).
Effectiveness: As shown in Table 5 column No Transfer, the Llama 8B model trained on the WDC-s-filtered and WDC-s-filtered-rel datasets outperforms the fine-tuned baseline by 4.73 and 3.18 F1 points, respectively. Both filtered approaches surpass the Llama 8B model trained on the WDC large training set, containing nearly 20,000 examples (see Table 1). In contrast, GPT-4o-mini does not exhibit similar improvements, as both filtered datasets perform below the zero-shot model.
Generalization: While filtration improves Llama 8B’s performance, it does not significantly impact in-domain generalization. The fine-tuned baseline captures 72% of dedicated performance, while filtered approaches capture 75% and 76% with relevancy filtration (see Table 5, In-Domain Transfer Gain column). Since GPT-4o-mini does not improve its performance with filtration in a no-transfer scenario, the same trend is observed in in-domain generalization, where this approach underperforms the fine-tuned baseline across all product datasets.
@@ 5.1. Training Set Filtration
Table 4. Impact of the filtration and generation methods on the size the training set.
@@
Table 5. Fine-tuning with example selection and generation. Deltas against fine-tuning on WDC-s
Model Train set No Transfer In-Domain Transfer Transfer Cross-Domain Transfer Transfer
@@
5.2. Example Generation
While the earlier approaches focused on discarding irrelevant examples, this section investigates using GPT-4o to generate additional training examples for fine-tuning. We test three methods for generating examples. All methods iterate over the WDC small training set and use the examples from this set as seeds derive additional training examples. The complete prompts used by the methods are provided in the project repository.
@@ 5.2. Example Generation
Detailed Generation Prompt: This prompt contains a comprehensive task description, including background information on entity matching and specific requirements for the generated examples. It instructs the model to produce examples from the same product category and to include similar matching challenges as the challenges found in the provided seed example. Additionally, the prompt explains key concepts like corner cases to deepen the model’s understanding of the task.
Demonstration-based Creation: Building on the second method, this approach includes six entity pairs as demonstrations into the prompt. These pairs are selected from the WDC small dataset based on their similarity to the seed example in the OpenAI embedding space.
@@
Combining Generation and Filtration: Given the mixed quality of the generated examples, we combine example generation with the filtering approach introduced above, further refined by relevancy filtering (syn.-filter-rel.). The resulting dataset sizes are presented in Table 4. Since the generation methods are intended to extend existing datasets, both training splits are combined with the unfiltered version of the WDC small dataset, and the counts in Table 4 reflect this combination.
Effectiveness: Table 5 column No Transfer shows that the Llama 8B model gains 3.35 F1 points without relevancy filtration and 4.85 F1 points with relevancy filtration. In contrast, GPT-4o-mini shows a 6.42 F1 point decrease compared to the fine-tuned baseline, underperforming zero-shot. Due to the high cost of fine-tuning the GPT series and the underperformance of this approach, we chose not to pursue fine-tuning with relevancy filtration.
@@
5.3. Error-based Example Selection
The idea of the error-based example selection (err.-sel.) approach is to guide the model to prevent errors by fine-tuning it with examples that are similar to entity pairs that are matched incorrectly by the model. For this, the Llama 8B model is initially trained on the standard 2,500 examples from the WDC small training dataset. After this initial training phase, the model is validated to identify the remaining errors. Additional training examples are then selected from the large WDC products dataset, as shown in Table 1, simulating additional labelling capacity. These examples are chosen based on their embedding similarity to the identified errors.
To maintain consistent training set sizes, each iteration starts with 2,500 examples from the WDC small dataset, adding 2,500 additional examples based on errors. The model is retrained on the expanded dataset for 5 epochs, and this process is repeated five times. Among all optimization rounds, the model with the highest F1 score on the validation set is selected as the best. Due to OpenAI’s fine-tuning limitations, this approach is only tested with the Llama series.
@@ 5.3. Error-based Example Selection
Generalization: This approach outperforms the generation method in a non-transfer scenario but underperforms in in-domain generalization, achieving 83% of dedicated performance compared to 72% for the fine-tuned baseline.
In cross-domain generalization, this approach shows a slight improvement over standard fine-tuning, with an average increase of 2.77 F1 points. However, it still lags behind the baseline model, which achieves, on average, 17.95 more F1 points when applied to the scholar domain.
6. Related work
Entity matching (Christen, 2012; Elmagarmid et al., 2007) has been researched for over 50 years (Fellegi and Sunter, 1969). While early techniques rely on hand-crafted matching rules, supervised machine learning techniques dominate the field today (Christophides et al., 2020; Barlaug and Gulla, 2021; Li et al., 2020; Peeters and Bizer, 2021; Yao et al., 2022; Zeakis et al., 2023; Wang et al., 2021). Narayan et al. (Narayan et al., 2022a) are the first to apply LLMs for entity matching using GPT-3. (Peeters and Bizer, 2023; Fan et al., 2023; Nananukul et al., 2024) extended Narayan’s work by testing a wider range of zero-shot and few-shot prompts on more recent models. While these approaches have shown strong performance, the prompt sensitivity of the LLMs in in-context learning settings has results in a costly and time-consuming search for the best prompt for each model/dataset combination (Lu et al., 2021; Margatina et al., 2023; Peeters and Bizer, 2024). Zhang et al. (Zhang et al., 2023) fine-tune LLMs using multi-task learning for handling various structured data-related tasks. Their set tasks include entity matching. More recently, Peeters et al. (Peeters and Bizer, 2024) experimented with fine-tuning the GPT3.5-Turbo-0613 model, achieving an F1 score of 88.34 on WDC Products. However, their results cannot be directly compared to the results in this paper as they are generated using only a subset of the WDC test set. Wadhwa et al. (Wadhwa et al., 2024) are the first to experiment with adding textual explanations to the training set used for fine-tuning LLMs. We extend this idea by augmenting the training set with different types of structured explanations. Wadhwa et al. note improvements, especially in out-of-domain scenarios. In contrast, our work shows that explanations primarily improve results in non-transfer and in-domain generalization settings.
7. Conclusion
This study demonstrates that standard fine-tuning provides significant performance improvements for smaller LLMs, as well as for certain larger models, such as GPT-4o. Small and large models benefit from augmenting the training set with structured explanations, which improve performance as well as generalization. The proposed example selection and generation methods only improve the performance of Llama 3.1 8B while decreasing the performance of GPT-4o Mini. As future work, we plan to refine the example selection and generation methods and develop strategies to improve cross-domain generalization.

### v2

@@ IV Dimension 1: Example Representation
This section investigates how different training example representations affect the performance and generalization capabilities of fine-tuned models. We compare three representations: 1) simple pairs as used in Section III, 2) pairs augmented by the LLM with textual explanations why the entities match or do not match, and 3) pairs augmented with structured explanations. The complete prompts for instructing the model to generate explanations as well as the augmented training sets are found in the project repository.
@@
Textual Explanations: We evaluate two types of textual explanations generated using GPT-4o-mini. The first approach presents the model with the prompt shown in Figure 2, followed by the label. Afterwards, the model is asked to explain the label. The model is given no guidance on the format of the explanation. This results in detailed explanations, having an average length of 293 tokens, which describe why the entity descriptions match or do not match. As a second approach, we re-run the technique proposed by Wadhwa et al. [18], which generates more concise explanations by including short explanations for other entity pairs as demonstrations in the prompt. The resulting explanations have an average length of 90 tokens. Figure 3 shows a training example containing an explanation in the AI part which was generated using this approach. The prompts used for generating these explanations, as well as the longer, more open-ended explanations, are available in the accompanying repository.
Figure 3: Training example containing a textual explanation.
@@
Structured Explanations: Our third approach is based on the assumption that LLMs can learn more from structured explanations compared to textual ones. Thus, building on the work of Peeters et al. [13], we instruct the model to generate explanations in a structured format that explicitly mention the attributes that were used for the matching decision, their importance for the decision, as well as the similarity of the attribute values. Figure 4 shows an example of a structured explanation that was generated by GPT-4o-mini. The complete prompt for the generation of the explanation is found in the project repository and is discussed in  [13]. The example shows that the LLM incorporates factors beyond string matching for calculating similarities, such as recognizing the relationship between Microsoft and Skype for Business.
Figure 4: Training example containing a Structured explanation.
@@
TABLE III: Fine-tuning results for different types of explanations (F1 scores, deltas to standard fine-tuning on WDC).
Model Train set No Transfer In-Domain Transfer Average Cross-Domain Transfer Average
@@
IV-A Effectiveness
We compare the effectiveness of different explanation approaches in a non-transfer scenario using the WDC Products dataset. The result of these experiments are found in the No Transfer column of Table III.
Smaller Models: Llama 8B outperforms standard fine-tuning across all evaluated explanation approaches, as shown in Table III, column No Transfer. Long textual explanations provide the smallest increase of 1.48 F1 points. The approach by Wadhwa et al. results in a 4.01-point improvement. Structured explanations further improve F1 scores by 0.93 over Wadhwa’s approach. GPT-4o-mini shows a different trend, where only structured explanations outperform the standard fine-tuning, increasing by 0.97 F1 points (Table III, WDC Products column). The other configurations underperform, with three out of five approaches performing worse than zero-shot (Table III, Non-transfer column).
Larger Model: Structured explanations are exclusively tested on larger models due to their consistent performance improvement in smaller models. While Llama 70B does not benefit from standard fine-tuning, structured explanations improve its performance by 1.50 F1 points over zero-shot (Table III, WDC column). In contrast, GPT-4o does not follow this trend. While structured explanations lead to a 1.60-point increase over zero-shot, standard fine-tuning results in a 5.50-point improvement, as shown in the GPT-4o rows in the WDC column of Table III. Therefore, GPT-4o does not significantly benefit from structured explanations in a non-transfer setting.
IV-B Generalization
@@
In-Domain Generalization: Standard fine-tuning results in Llama 8B achieving 72% of the dedicated models’ performance on product datasets (Table III, In-Domain Average Gain column). The Wadhwa et al. approach and longer textual explanations underperform, achieving 55% and 51%, respectively. However, structured explanations surpass the fine-tuned baseline, with the structured and no importance approaches achieving 93% and 91% of source model performance (Table III, In-Domain Average Gain column). GPT-4o-mini continues the trend where 4 out of 5 explanation approaches underperform the fine-tuned baseline. However, structured explanations outperform the baseline, achieving 23% of the source models’ performance compared to 13% by the fine-tuned baseline (Table III, In-Domain Average Gain column). Llama 70B benefits from structured explanations, outperforming both the fine-tuned baseline and zero-shot in in-domain generalization. It shows the highest generalization performance on two out of three product datasets (Table III, Llama 70B structured columns A-B and W-A). This trend however does not extend to GPT-4o as structured explanations underperform in this instance. Llama 70B fine-tuned with structured explanations outperforms the other approaches. Specifically, it achieves the highest generalization gain on two out of three product datasets (Table III, Llama 70B structured, columns A-B and W-A). However, this improvement does not extend to GPT-4o, where structured explanations negatively impact generalization performance.
Cross-Domain Generalization: Llama 8B struggles to generalize effectively in cross-domain settings, with all approaches underperforming compared to zero-shot performance. In contrast, GPT-4o-mini shows improvements across all explanation approaches, however non-rival zero-shot performance. In conclusion, structured explanations significantly enhance in-domain generalization, while cross-domain generalization remains limited.
V Dimension 2: Example Selection and Generation
@@
V-A Training Set Filtration
The filtration approach is based on the assumption that training data quality outweighs quantity and it is thus beneficial to remove potentially misleading examples from the training set. This is in accordance with the documentation of OpenAI who recommend smaller training sizes in favor of presenting examples of missbehaviour by the model [^6] https://platform.openai.com/docs/guides/fine-tuning.
Error-based Filtering: This approach attempts to filter out mislabeled training examples from WDC training set. For this purpose, we use GPT-4o-mini and discard all examples that are incorrectly labeled by the model. This approach may discard some correctly labeled examples. However, it is hypothesized that overall dataset quality improves. The resulting training set WDC-filtered has a size of 2006 examples, compared to 2500 examples of the original WDC-small training set (see Table IV).
Relevancy-based Filtering: The second approach involves filtering for relevance, operating on the assumption that fewer, more meaningful examples lead to better outcomes. For example, comparing a hard drive and a TV might pass initial filtration but offer limited value due to their low complexity. By removing such examples, computational demands are certainly reduced, while training results may potentially be enhanced. This approach prompts GPT-4o to select only “interesting” training examples, a term we intentionally leave undefined, allowing the model to interpret it. The model appears to define “interesting” as pairs that share many attributes and are thus highly similar (corner cases). Pairs deemed irrelevant by the model are discarded. The resulting WDC-filtered-rel training set has a total size of 608 examples (see Table IV).
Effectiveness: As shown in Table V column No Transfer, the Llama 8B model trained on the WDC-s-filtered and WDC-s-filtered-rel datasets outperforms the fine-tuned baseline by 4.73 and 3.18 F1 points, respectively. Both filtered approaches surpass the Llama 8B model trained on the WDC large training set, containing nearly 20,000 examples (see Table I). In contrast, GPT-4o-mini does not exhibit similar improvements, as both filtered datasets perform below the zero-shot model.
Generalization: While filtration improves Llama 8B’s performance, it does not significantly impact in-domain generalization. The fine-tuned baseline captures 72% of dedicated performance, while filtered approaches capture 75% and 76% with relevancy filtration (see Table V, In-Domain Average Gain column). Since GPT-4o-mini does not improve its performance with filtration in a no-transfer scenario, the same trend is observed in in-domain generalization, where this approach underperforms the fine-tuned baseline across all product datasets.
@@ V-A Training Set Filtration
TABLE IV: Training set sizes after filtering and generation steps.
@@
TABLE V: Fine-tuning with example selection and generation. Deltas against fine-tuning on WDC-small.
Model Train set No Transfer In-Domain Transfer Average Cross-Domain Transfer Average
@@
V-B Example Generation
While the earlier approaches focused on discarding irrelevant examples, this section investigates using GPT-4o to generate additional training examples for fine-tuning. We test three methods for generating examples. All methods iterate over the WDC small training set and use the examples from this set as seeds to derive additional training examples. The complete prompts used by the methods are provided in the project repository.
@@ V-B Example Generation
Detailed Generation Prompt: This prompt contains a comprehensive task description, including background information on entity matching and specific requirements for the generated examples. It instructs the model to produce examples from the same product category and to include similar matching challenges as the challenges found in the provided seed example. In addition, the prompt explains key concepts such as corner cases to deepen the model’s understanding of the task.
Demonstration-based Creation: Building on the second method, this approach includes six entity pairs as demonstrations in the prompt. These pairs are selected from the WDC small dataset based on their similarity to the seed example in the OpenAI embedding space.
@@
Combining Generation and Filtration: Given the mixed quality of the generated examples, we combine example generation with the filtering approach introduced above, further refined by relevancy filtering (syn.-filter-rel.). The resulting dataset sizes are presented in Table IV. Since the generation methods are intended to extend existing datasets, both training splits are combined with the unfiltered version of the WDC small dataset (Table IV).
Effectiveness: Table V column No Transfer shows that the Llama 8B model gains 3.35 F1 points without relevancy filtration and 4.85 F1 points with relevancy filtration. In contrast, GPT-4o-mini shows a 6.42 F1 point decrease compared to the fine-tuned baseline, underperforming zero-shot. Due to the high cost of fine-tuning the GPT series and the underperformance of this approach, we chose not to pursue fine-tuning with relevancy filtration.
@@
V-C Error-based Example Selection
The idea of the error-based example selection (err.-sel.) approach is to guide the model to prevent errors by fine-tuning it with examples that are similar to entity pairs that are matched incorrectly by the model. For this, the Llama 8B model is initially trained on the standard 2,500 examples from the WDC small training dataset. After this initial training phase, the model is validated to identify the remaining errors. Additional training examples are then selected from the large WDC products dataset, as shown in Table I, simulating additional labeling capacity. These examples are chosen based on their embedding similarity to the identified errors.
Each iteration starts with 2,500 examples from the WDC small dataset, adding 2,500 more based on errors. The model is retrained for 5 epochs, repeating this process five times. The best model is selected based on the highest F1 score on the validation set. Due to OpenAI’s fine-tuning limitations and high computational cost, this approach is tested only with Llama 8B.
@@ V-C Error-based Example Selection
Generalization: While this approach yields improved dedicated performance, generalization does not improve beyond that of example generation (83% average gain). Interestingly, this approach results in severe underperformance in a cross-domain transfer scenario, with the worst results out of all tested approaches.
VI Cost Analysis
This chapter presents a cost analysis for hosted LLMs. Table VI lists the costs and the number of tokens used by GPT-4o-mini and GPT-4o for the WDC Products dataset across three scenarios: zero-shot, standard fine-tuning, and structured explanation fine-tuning.
TABLE VI: Cost comparison GPT-4o-mini/4o on WDC-small.
Zero-shot Fine-tuned Standard Fine-tuned Structured
Metric 4o-mini 4o 4o-mini 4o 4o-mini 4o
F1 score 81.61 81.60 83.41 87.10 84.38 83.20
Fine-tuning (2,500 examples)
Training tokens 0 0 1,841,460 1,841,460 5,750,330 5,750,330
Cost per example 0 0 0.22¢ 1.84¢ 0.69¢ 5.75¢
Total fine-tuning cost 0 0 \$5.52 \$46.04 \$17.25 \$143.76
Inference (4,500 examples)
Total Input tokens 338,735 338,735 338,735 338,735 338,735 338,735
Total Output tokens 4,500 4,626 4,500 4,500 14,758 14,683
Mean Token Count 76.27 76.30 76.27 76.27 78.55 78.54
Total Token Count 343,235 343,361 343,235 343,235 353,493 353,418
Total Inference Cost \$0.05 \$0.89 \$0.11 \$1.34 \$0.12 \$1.49
At the time of our experiments, GPT-4o-mini costs \$0.15 (GPT-4o: \$2.50) per million input tokens and \$0.60 (GPT-4o: \$10.00) per million output tokens in a zero-shot setting. Fine-tuning is billed at \$3 (GPT-4o: \$25.00) per million training tokens. Additionally, inference on the resulting fine-tuned model doubles the token price to \$0.30 (GPT-4o: \$3.75) per million input tokens and \$1.20 (GPT-4o: \$15.00) per million output tokens. [^7] https://openai.com/api/pricing/ (as of January 2025) This pricing structure generally makes fine-tuning more expensive than zero-shot usage. As shown in Table VI, zero-shot inference on GPT-4o-mini and GPT-4o yields comparable F1 scores (81.61 vs. 81.60), but GPT-4o is significantly more expensive at \$0.89 for inference compared to \$0.05 for GPT-4o-mini. Standard fine-tuning improves GPT-4o-mini’s F1 to 83.41 at a \$5.52 fine-tuning cost, while GPT-4o jumps to 87.10 at a fine-tuning cost of \$46.04. Structured fine-tuning increases GPT-4o-mini’s performance further (F1=84.38) for \$17.25, whereas GPT-4o drops slightly (F1=83.20) for \$143.76 in fine-tuning costs. Thus, while GPT-4o obtains a larger gain in standard fine-tuning, it also incurs substantially higher training and inference expenses. By contrast, GPT-4o-mini demonstrates that more elaborate techniques (e.g., structured fine-tuning) can remain cost-effective. Whether GPT-4o’s higher scores justify the tenfold cost depends on requirements of the specific use case. Ultimately, these findings highlight that a model’s raw performance gains must be weighed against sharply escalating token and fine-tuning costs, particularly for hosted large-scale models. The cost advantage of models without fine-tuning likely holds only in a hosted scenario; in self-hosted settings, there should not be a difference in the inference cost between a model without fine-tuning and with fine-tuning.
VII Related work
Entity matching [2, 3] has been studied extensively over the last 50 years [19]. Early methods relied on manually defined matching rules, while the field is now dominated by supervised machine learning techniques [4, 1, 6, 7, 20, 9, 21]. A prominent system leveraging pre-trained language models (PLMs), such as BERT, is Ditto[6]. PLM-based approaches typically require fewer computational resources, due to their smaller size, compared to LLM-based matchers [13].
Entity Matching using LLMs: Narayan et al.[11] were the first to experiment with employing LLMs for entity matching. Subsequent studies[12, 14, 22] expanded this approach by evaluating further zero-shot and few-shot prompts on more recent models. Peeters, et al. [13] investigate the effect of standard fine-tuning on the effectiveness and generalization performance of GPT4o-mini, Llama2-70B, and Llama3.1-70B. We build upon this work and experiment with varying the representation of training examples and use different example selection and generation methods. The fine-tuning results from  [13] for specific datasets are not directly comparable to the results that we present in Chapter III, as they subsample the test sets in order to save resources, while we use the complete test sets.
Fine-tuning using Explanations: Zhang et al.[15] explored fine-tuning LLMs via multi-task learning for multiple data integration tasks, including entity matching. They provide the models with reasoning traces in the training data, going beyond standard fine-tuning. Wadhwa et al.[18] expanded upon this by incorporating textual explanations into the training process while focusing solely on entity matching. In contrast, we include structured explanations in the training data, hypothesizing that the consistent structure makes it easier for the model to learn from these explanations compared to purely textual approaches, such as [15, 18].
Generalization Performance: While earlier work [8, 23] showed that PLMs, such as BERT, exhibit poor generalization for entity matching, AnyMatch by Zhang et al.[24] demonstrates that smaller models can generalize in a leave-one-dataset-out evaluation setup. They also show that using attribute-level transformations for training data augmentation can increase the generalization performance. Notably, their use of GPT-2, a decoder-based architecture, differs from the encoder-based architectures studied in[8]. In contrast, our method explicitly generates novel synthetic examples, exploring a different strategy to further improve model robustness and generalization. In addition, we differentiate between in-domain and cross-domain generalization, taking into account the varying levels of domain shift. Li, et al. [25] investigate the ability of LLMs to generalize to unseen table tasks in a multi-task fine-tuning setup.
VIII Conclusion
This paper shows that by fine-tuning small LLMs for entity matching, it is possible to reach a performance comparable to that of large models. This significantly reduces both costs and the CO2 footprint associated with using LLMs for entity matching. Fine-tuning does not harm the model’s ability to generalize to unseen datasets, while fine-tuning with structured explanations even improves this capability. In future work, we plan to refine the example selection and generation methods and explore strategies to improve cross-domain generalization.